Circular No. 8325

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 H1 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to be diffuse by at least two different CCD observing groups, including M. Kočer at Kleť (1.06-m KLENOT telescope on Apr. 16.8 UT) and A. Gilmore and P. Kilmartin at Mt. John (0.6-m reflector on Apr. 18.4).

2004 UT  α_{2000}  δ_{2000}  Mag.
Apr. 16.13007  7^{h}42^{m}04.^{s}72  +0^{o}13′25′′.2  19.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-H26.

T = 2004 Jan. 13.329 TT  ω = 288.^{o}723
Ω = 119.801 }2000.0
q = 2.04920 AU  i = 141.155

GRAVITATIONAL MICROLENSING EVENT G. Richards, D. Johnston, and J. Hennawi, Princeton University Ob-

servatory, report that an emission-line enhancement has been observed for the second time in component A of the wide-separation, quadruply imaged quasar SDSS J100434.82+411239.5 (component A, α = 10^{h}04^{m}34^{s}.82, δ = +41^{o}12′39′′.5, equinox 2000.0; magnitude i = 18.4; Inada et al. 2003, Nature 426, 810). Specifically, the blue wings of several high ionization emission lines were enhanced in component A on Mar. 26 and Apr. 10 with respect to observations made in 2003 Nov./Dec., based on data taken with the ARC 3.5-m telescope at Apache Point Observatory. This second epoch of enhancement follows that of 2003 May, confirming the time-dependent nature of this possible microlensing event as reported by Richards et al.

(2004, Ap.J., in press; http://xxx.lanl.gov/abs/astro-ph/?0402345).

No enhancement has been seen in component B (position end figures 34^{s}.94, 43′′.0), despite the short A-B time delay (< 30 days) and good time sampling. Follow-up observations of components A and B are needed to confirm the microlensing nature of the event and to probe both the microlens distribution and the structure of the broad emission-line region. Observers may contact the first author at gtr@astro.princeton.edu for more information or to coordinate observations.

2004 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green